A glass block of length 18 cm and refractive index 1.5 contains a small bubble. Viewed from one side the bubble appears to be 8 cm from this side. How far in the block will it appear to be when viewed from the opposite side?
1
Expert's answer
2020-09-21T08:31:35-0400

Refractive index = real depth/apparent depth

Let the bubble is at distance x from the side F1 of the glass block.

"1.5 = \frac{x}{8}"

"x = 12\; cm"

The bubble is at distance ("18-12 = 6 \;cm" ) opposite side F2 of the glass block.

Let the y is apparent depth when viewed from the side F2 of the glass block.

"1.5 = \frac{6}{y}"

"y = 4 \;cm"

Answer: 4 cm